OpenStudy (anonymous):

The function f is such that f\[ f(x) = 3(x+2)^3 -5~~~for ~~x \ge 0\](i) Obtain an expression for \[f'(x)\] and hence explain why f is an increasing function.

OpenStudy (anonymous):

@dpaInc

OpenStudy (hoblos):

f'(x) = 9(x+2)² since x>0 x+2 >0 9(x+2)² >0 so f'(x) >0 thus f is an increasing function

OpenStudy (anonymous):

How did you get f'(x)?

OpenStudy (hoblos):

the formula says: \[(u ^{n})' = n(u ^{n-1})u'\] so [ 3(x+2)^3 -5]'= [3(x+2)^3]' - (5)' = 3(3)[(x+2)^(3-1)] * (x+2)' - 0 = 9(x+2)² (1) = 9(x+2)²

OpenStudy (anonymous):

you good?

OpenStudy (anonymous):

Yup! :D